# Select to Know: An Internal-External Knowledge Self-Selection Framework for Domain-Specific Question Answering

### v1

@@ 1 Introduction
  • Experiments across the medicine, law, and finance demonstrate that S2K matches pretrained LLMs with significantly lower training cost.
@@ 3.1.2 Internal-External Fusion Knowledge
$t_{i}^{F}=\begin{cases}t_{i}^{I},&\text{if}P_{M}(t_{i}^{I}\mid Q,A^{F}_{<i})>P_{M}(t_{i}^{E}\mid Q,D,A^{F}_{<i})\\ t_{i}^{E},&\text{otherwise}\end{cases}$} (2)
Here, $A_{<i}^{F}=\{t_{0}^{F},t_{1}^{F},t_{2}^{F},…,t_{i-1}^{F}\}$, which ensures two key properties: (1) the final answer fused from internal and external knowledge remains coherent and readable, and (2) the only difference between the two inference settings is whether the external document $D$ is provided.
@@
$t_{i:i+W}^{F}=\begin{cases}t_{i:i+W}^{I},&\text{if }\begin{aligned} &\frac{1}{W}\sum_{j=0}^{W-1}\log P_{M}(t_{i+j}^{I}\mid Q,A^{F}_{<i})\geq\\ &\frac{1}{W}\sum_{j=0}^{W-1}\log P_{M}(t_{i+j}^{E}\mid Q,D,A^{F}_{<i})\\ &+C\end{aligned}\\ t_{i:i+W}^{E},&\text{otherwise}\end{cases}$ (3)
@@ 3.2 Internal–External Knowledge Fusion Training
$\omega_{t}=(1-\text{correct}_{t})+\text{correct}_{t}\cdot\frac{H_{t}}{\log V}$ (5)
@@
$\mathcal{L}=\frac{1}{N}\sum_{t=1}^{T}\omega_{t}\cdot\text{NLL}_{t}$ (6)
@@ 3.3 Reasoning-Enhanced Training
After acquiring domain knowledge, we apply GRPO, a critic-free reinforcement learning method, to improve the reasoning capabilities of the LLM. Following raw GRPO, we design an accuracy reward and a format reward. The accuracy reward ($R_{\text{acc}}$) has two cases: +5 for a fully correct answer and 0 for an incorrect one. The format reward ($R_{\text{fmt}}$) includes three cases: +1 for strictly following the "<think>…</think>…ANSWER" format, 0 for a general formatting error, and –0.5 if "ANSWER" is generated multiple times, which indicating a potential reward-hacking behavior, where the model outputs multiple candidate answers to maximize reward. The final reward is the sum of both: $R=R_{\text{acc}}+R_{\text{fmt}}$.
@@ 4.3.3 Relevance-based sampling of Reasoning Data Generation
Table 3: Comparison of reward schemes. While Acc means Accuracy reward, Fmt means format reward and EA means extra-answer penalty.
@@ 4.3.4 Reward Function Analysis
We use GRPO with accuracy and format rewards to boost QA performance in real-world, domain-specific settings. We compare three reward schemes: (1) Answer Only: binary reward for answer correctness; (2) Answer + Format: combined reward for correctness and formatting; and (3) Enhanced Answer + Format: combined reward with stronger Answer incentives.
@@
Std. SFT & Ext. Data 33.5 36.8 68.7
Sel. SFT & Ext. Data 34.2 37.9 73.1
Sel. SFT & Fus. Data 35.1 39.6 75.9
Only GRPO 36.1 36.4 61.4
S2K 38.6 43.4 77.1
Table 4: Ablation study.   Internal–External Fusion fine-grained ablation;   End to End ablation; (Abbreviations: Std. SFT=Standard SFT; Sel. SFT=Selective SFT; Ext. Data=External Training Data; Fus. DT=Fusion Training Data)
@@
4.4 Ablation Study
To further validate the contribution of each component in S2K, we conduct a detailed ablation study covering internal-external knowledge fusion, reinforcement learning, and end-to-end training. As shown in Table 4, during the first-stage training, our proposed Selective SFT (Section 3.2) outperforms standard SFT, and the fusion of internal and external knowledge (Section 3.1.2) leads to better performance than using external knowledge alone, demonstrating the effectiveness of both components. Furthermore, compared to the final model trained with the full two-stage pipeline (Avg@5: 38.7), models trained with only Selective SFT (35.2) or only GRPO (36.5) exhibit inferior performance, highlighting the importance and effectiveness of our overall training strategy.
4.5 Case Study
In this section, we present a real case in Table 5, to demonstrate how our fusion mechanism works. The original document describes the functional differences between M1 and M2 macrophages in arginine metabolism, while the external, though containing some relevant facts, introduces notable knowledge errors. Our fusion answer successfully identifies and corrects these errors while retaining complementary details from the external source, resulting in a more complete and accurate answer.
@@ 5 Related Work
Continued Training Domain Adaptation: Continued training Labrak et al. (2024); Qiu et al. (2024); Zhang et al. (2025); Mecklenburg et al. (2024) aims to inject domain-specific knowledge into LLMs to compensate for their lack of specialized expertise. This strategy can be broadly divided into two main approaches: pre-training Qiu et al. (2024); Shu et al. (2024); Li et al. (2024); Chen et al. (2023) adaptation, which fine-tunes LLMs on domain-specific corpora to help them internalize expert knowledge (e.g., BioMistral Labrak et al. (2024)); and post-training Zhang et al. (2025); Mecklenburg et al. (2024), which involves fine-tuning LLMs using QA pairs derived from domain knowledge. However, continued training often encounters hurdles in effectively enabling LLMs to extract the acquired knowledge during the inference phase  Zhang et al. (2025); Ibrahim et al. ; Ovadia et al. (2024).
@@ 6 Conclusion
To address challenges in DSQA, we propose S2K, an efficient framework designed to enhance the performance of LLMs in long-tail domains. In vertical domains where no readily available QA datasets exist, S2K enables effective transfer and generalization of QA capabilities using only raw corpora. Experiments across multiple representative vertical domain results demonstrate its effectiveness.
@@ A.2.2 Hyperparameter
In addition, Table 8 provides the detailed hyperparameter settings used in the GRPO stage. For fair comparison, the reinforcement learning baselines are configured with the same hyperparameters.
@@ A.2.3 Metric
$\text{Cons@5}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\text{major}(a_{i1},\dots,a_{i5})=a_{i}^{\text{gold}}\right)$ (8)
@@
$\text{Pass@5}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\sum_{j=1}^{5}y_{ij}\geq 1\right)$ (9)
@@ A.5 Prompts
We generate meta knowledge and reasoning data by prompting DeepSeek-v3 and GPT-4o. This section provides detailed prompt templates. Table 11 presents the prompt used for meta knowledge generation, while Tables 12 to 14 show the prompts used for generating inductive, deductive, and case-based reasoning data, respectively.

### v2

@@ 1 Introduction
  • Experiments across the medicine, law, and finance demonstrate that S2K matches pretrained LLMs with significantly lower cost.
@@ 3.1.2 Internal-External Fusion Knowledge
$t_{i}^{F}=\begin{cases}t_{i}^{I},&\text{if}P_{M}(t_{i}^{I}\mid Q,A^{F}_{<i})>P_{M}(t_{i}^{E}\mid Q,D,A^{F}_{<i})\\ t_{i}^{E},&\text{otherwise}\end{cases}$
Here, $t_{i}^{I}$ and $t_{i}^{E}$ refer to the model’s token-level predictions under internal-only knowledge and external knowledge, respectively. $A_{<i}^{F}=\{t_{0}^{F},t_{1}^{F},t_{2}^{F},\dots,t_{i-1}^{F}\}$, which ensures two key properties: (1) the final answer fused from internal and external knowledge remains coherent and readable, and (2) the only difference between the two inference settings is whether the external document $D$ is provided.
@@
$t_{i:i+W}^{F}=\begin{cases}t_{i:i+W}^{I},&\text{if }\begin{aligned} &\frac{1}{W}\sum_{j=0}^{W-1}\log P_{M}(t_{i+j}^{I}\mid Q,A^{F}_{<i})\geq\\ &\frac{1}{W}\sum_{j=0}^{W-1}\log P_{M}(t_{i+j}^{E}\mid Q,D,A^{F}_{<i})\\ &+C\end{aligned}\\ t_{i:i+W}^{E},&\text{otherwise}\end{cases}$
@@ 3.2 Internal–External Knowledge Fusion Training
$\omega_{t}=(1-\text{correct}_{t})+\text{correct}_{t}\cdot\frac{H_{t}}{\log V}$
@@
$\mathcal{L}=\frac{1}{N}\sum_{t=1}^{T}\omega_{t}\cdot\text{NLL}_{t}$
@@ 3.3 Reasoning-Enhanced Training
After acquiring domain knowledge, we apply GRPO, a critic-free reinforcement learning method, to improve the reasoning capabilities of the LLM. We design an accuracy reward and a format reward. The accuracy reward ($R_{\text{acc}}$) has two cases: +5 for a fully correct answer and 0 for an incorrect one. The format reward ($R_{\text{fmt}}$) includes three cases: +1 for strictly following the "<think>…</think>…ANSWER" format, 0 for a formatting error, and –0.5 if "ANSWER" is generated multiple times, indicating potential reward hacking, where the model outputs multiple candidate answers to maximize reward. The final reward is the sum of both: $R=R_{\text{acc}}+R_{\text{fmt}}$.
@@ 4.3 Analysis Experiments
Model MedQA JECQA FinanceIQ
Avg@5 Cons@5 Pass@5 Avg@5 Cons@5 Pass@5 Avg@5 Cons@5 Pass@5
LLaMA3.1-8B 22.0 22.5 70.9 8.08 18.3 26.7 12.7 7.6 47.7
   +Sel.SFT 27.3 31.3 78.9 9.76 20.3 31.6 21.5 23.4 70.0
   +GRPO 29.2 32.4 81.0 20.0 23.0 46.8 23.0 24.7 73.0
Table 2: Evaluation of S2K on a different model architecture. Applying Selective SFT and GRPO on LLaMA3.1-8B consistently improves performance across MedQA, JECQA, and FinanceIQ, showing that our framework generalizes beyond the Qwen family of models.
Model Avg@5 Cons@5 Pass@5
Qwen2.5-1.5B 8.9 9.7 36.9
   +S2K 17.0 15.1 60.3
Table 3: Evaluation of S2K on a smaller model(Qwen2.5-1.5B) on the FinanceIQ. Despite the limited parameter size, applying our framework leads to substantial improvements across all metrics, demonstrating the scalability of S2K.
@@ 4.3.2 Window Width $W$ in Knowledge Fusion
4.3.3 Robustness Across Model Architectures and Sizes
To evaluate the robustness of S2K across different model architectures and parameter scales, we conduct experiments on LLaMA3.1-8B Dubey et al. (2024) and Qwen2.5-1.5B. As shown in Table 2, applying Selective SFT on LLaMA3.1-8B already yields consistent improvements across MedQA, JECQA, and FinanceIQ. Further incorporating GRPO leads to substantial gains, demonstrating the effectiveness of S2K.
In addition, Table 3 shows that S2K also benefits smaller models such as Qwen2.5-1.5B. Even with limited model capacity, our framework significantly enhances performance on all three metrics, suggesting that S2K is general and scalable across different architectures and model sizes.
@@ 4.3.4 Relevance-based sampling of Reasoning Data Generation
Table 5: Comparison of reward schemes. While Acc means Accuracy reward, Fmt means format reward and EA means extra‐answer penalty.
@@ 4.3.5 Reward Function Analysis
We use GRPO with accuracy and format rewards to boost QA performance in real-world, domain‐specific settings. We compare three reward schemes: (1) Answer Only: binary reward for answer correctness; (2) Answer + Format: combined reward for correctness and formatting; and (3) Enhanced Answer + Format: combined reward with stronger Answer incentives.
@@
Only Sel.SFT $35.1_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+1.6}}}$ $39.6_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+1.3}}}$ $75.9_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+8.3}}}$
Only GRPO $36.1_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+2.6}}}$ $36.4_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.9}}}$ $61.4_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-6.2}}}$
[…]
Table 6: Ablation study on training stages. We compare the raw LLM, applying only Selective SFT, only GRPO, and their combination. The results show that Selective SFT and GRPO are both beneficial, while their combination yields the best overall performance.
Method Data Avg@5 Cons@5 Pass@5
- - $33.5$ $38.3$ $67.6$
Std.SFT External $33.5_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0}}}$ $36.8_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.5}}}$ $68.7_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+1.1}}}$
Sel.SFT External $34.2_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.7}}}$ $37.9_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.4}}}$ $73.1_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+5.5}}}$
[…]
Table 7: Ablation study on SFT. We examine standard SFT and Selective SFT under different data settings. Selective SFT consistently outperforms standard SFT, and incorporating fusion knowledge further enhances all metrics.
4.4 Ablation Study
To provide a clearer understanding of the contribution of each component in S2K, we report ablation studies from three complementary perspectives.
First, Table 6 presents the effect of different training stages. We observe that applying only Selective SFT leads to notable improvements over the raw LLM, while GRPO alone slightly improves Avg@5 but causes degradation in Cons@5 and Pass@5. In contrast, combining Selective SFT with GRPO yields the best results across all metrics, highlighting the effectiveness of our two-stage training pipeline.
Second, Table 7 focuses on the SFT stage. Compared with standard SFT using external knowledge, Selective SFT achieves higher accuracy, especially in terms of Pass@5. Moreover, when incorporating our proposed internal–external fusion knowledge, Selective SFT further boosts performance across all metrics.
Finally, Table 8 compares our fusion-based training data with data obtained by directly distilling R1. While both strategies improve upon the raw LLM, the fusion data leads to consistently larger gains across Avg@5, Cons@5, and Pass@5. These results demonstrate that our data fusion approach provides higher-quality supervision than direct distillation, and is a key factor behind the effectiveness of S2K in domain-specific QA.
4.5 Case Study
Method Avg@5 Cons@5 Pass@5
Raw LLM $18.0$ $17.7$ $62.2$
+R1-distill Data $20.0_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+2.0}}}$ $20.3_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+2.6}}}$ $66.7_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+4.5}}}$
+Fusion Data $22.2_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+4.2}}}$ $24.0_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+6.3}}}$ $71.6_{\text{{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+9.4}}}$
Table 8: Comparison of training data quality on the FinanceIQ dataset. Both R1-distilled data and our fusion data improve over the raw LLM, but the fusion data provides consistently larger gains across all metrics, demonstrating the superiority of our data fusion strategy.
In this section, we present a real case in Table 9, to demonstrate how our fusion mechanism works. The example focuses on the functional differences between M1 and M2 macrophages in arginine metabolism. The original document accurately captures their metabolic roles, whereas the external source, despite providing partially relevant details, introduces factual errors by incorrectly linking M1 to Th2 and M2 to Th1 responses. Our fusion answer not only corrects these mistakes but also retains complementary information from the external source, resulting in a response that is both more accurate and more comprehensive.
@@ 5 Related Work
Continued Training Domain Adaptation: Continued training Labrak et al. (2024); Qiu et al. (2024); Zhang et al. (2025); Mecklenburg et al. (2024) aims to inject domain-specific knowledge into LLMs to compensate for their lack of specialized expertise. This strategy can be broadly divided into two main approaches: pre-training Qiu et al. (2024); Shu et al. (2024); Li et al. (2024); Chen et al. (2023) adaptation, which fine-tunes LLMs on domain-specific corpora to help them internalize expert knowledge (e.g., BioMistral Labrak et al. (2024)); and post-training Zhang et al. (2025); Mecklenburg et al. (2024); Tian et al. (2023), which involves fine-tuning LLMs using QA pairs derived from domain knowledge. However, continued training often encounters hurdles in effectively enabling LLMs to extract the acquired knowledge during the inference phase  Zhang et al. (2025); Ibrahim et al. ; Ovadia et al. (2024). Recent studies further highlight the importance of data diversity for improving generalization during fine-tuning Song et al. (2024), and propose knowledge-aware fine-tuning strategies to explicitly inject and utilize external knowledge, thereby mitigating hallucinations Lyu et al. (2024).
@@ 6 Conclusion
To address challenges in DSQA, we propose S2K, an efficient framework designed to enhance LLM performance in long-tail domains. In vertical domains where no readily available QA datasets exist, S2K enables effective transfer and generalization of QA capabilities using only raw corpora. Experiments across multiple representative vertical domains demonstrate its effectiveness. In addition to strong accuracy, S2K achieves comparable performance to domain-pretrained models at significantly lower cost.
@@ A.2.2 Hyperparameter
In addition, Table 12 provides the detailed hyperparameter settings used in the GRPO stage. Specifically, the global batch size refers to the number of total training samples in one optimization step (set to 1 in our case due to resource constraints), while the rollout batch size denotes the number of rollouts per sample (set to 8), balancing computational efficiency and diversity in policy updates. For fair comparison, the reinforcement learning baselines are configured with the same hyperparameters.
@@ A.2.3 Metric
$\text{Cons@5}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\text{major}(a_{i1},\dots,a_{i5})=a_{i}^{\text{gold}}\right)$
@@
$\text{Pass@5}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\sum_{j=1}^{5}y_{ij}\geq 1\right)$
@@ A.5 Prompts
We generate meta knowledge and reasoning data by prompting DeepSeek-v3 and GPT-4o. This section provides detailed prompt templates. Table 15 presents the prompt used for meta knowledge generation, while Tables 16 to 18 show the prompts used for generating inductive, deductive, and case-based reasoning data, respectively.